Make New Entry,
Make Followup Entry
User name mercado
Log entry time 15:26:04 on August29,2009
Entry number 286838
Cryo Alarms going off. Dave Meekins was called in. Calling for access to try to fix problem.
PID controller seems to be not functioning, and auxilliary heater controller is unable to produce full power (~1kw) to keep target warm. Beam used at 20 uA to control temp, which stabilized. Later, 750W from aux heater is suffient to stay around 16.4K.
Meekins to the hall in controlled access to diagnose heater power supply... hopefully cryo flows don't go up before heater is recovered!